OpenStudy (anonymous):

for all a>1, the expression 3a^4/3a^6 equals? how do i do this question?

OpenStudy (mathteacher1729):

They just want you to simplify: \[\huge \frac{3a^4}{3^6}=\cancel{\frac{3}{3}}a^{4-6}=a^{-2}=\frac{1}{a^2}\]

OpenStudy (anonymous):

could you give me a similar equation so i can see if i can do it?

OpenStudy (anonymous):

Here's a similar-level exercise for you: simplify \(\dfrac{15b^3}{3b^2}\).

OpenStudy (anonymous):

1/5b

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Close, but not quite. Try again :)

OpenStudy (anonymous):

Hint: \[\frac{15b^3}{3b^2}=\frac{5\cdot 3 \cdot b\cdot b \cdot b}{3 \cdot b \cdot b}=?\]

OpenStudy (anonymous):

then the other answer would have been 5b

OpenStudy (anonymous):

Correct. The terms that "survive" the cancellation process are in the numerator, not the denominator. \[\frac{15b^3}{3b^2}=\frac{5\cdot \cancel 3 \cdot \cancel b\cdot \cancel b \cdot b}{\cancel 3 \cdot \cancel b \cdot \cancel b}=5b\]

OpenStudy (anonymous):

oh ok thank you muches... i get it now

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

:)